# Ego-to-Exo: Interfacing Third Person Visuals from Egocentric Views in Real-time for Improved ROV Teleoperation

### v2

@@ III-B Generating Exocentric Perspective
Here, $\lambda_{2}$ is the scaling factor for the ROV model. Like all monocular SLAM-based systems [47, 48], our proposed mapping and projection method is up to scale. While the scale can be resolved with additional sensor fusion, the augmented visuals of Eq. 3 are sufficient for teleoperation.
@@ IV-A Implementation Details
The proposed framework is implemented in ROS Noetic on an Ubuntu 20.04 environment. A ROS node for ORB-SLAM3 is integrated as the monocular SLAM backbone. The buffer queue size $n$ is set to $100$ and the frame separation threshold is set to $0.001$ unit. The ROV point clouds are generated by sampling 3D mesh models of BlueROV2 and TurtleBot4; $10,000$ points are sampled for each model. The same ROS packages are used on both robots with domain-specific empirical tuning for the $\lambda$ parameters.
@@ IV-B Proof of Concept: 2D Indoor Navigation
Experimental setup. The proof-of-concept experiments are conducted with TurtleBot4, a 2D ground robot that can be teleoperated with egocentric views from its front-facing monocular camera. It has only two degrees of freedom (DOF) for linear and angular velocity - which simplifies the motion kinematics for tracking its instantaneous position and orientation. We teleoperate it to collect monocular visual data in office, laboratory, and hallway scenarios. The experiments are designed to validate the proposed algorithm by evaluating ground plane estimation and reprojection errors.
@@ IV-C Field Deployment: 3D Underwater Cave Exploration
Experimental setup. We extend our experiments to underwater cave exploration scenarios, where the ROV performs full 6-DOF motions. While the $roll$ motion is limited in the standard BlueROV2s, we consider all 6-DOF for teleoperation with the buoyancy change and pressure imbalance caused by water flow at the cave openings. For remote teleoperation, we consider the scenarios where human operators maneuver an underwater ROV from the surface by following the caveline and other navigation markers as a guide [49]. The mission objective is to navigate the ROV $75$-$300$ feet deep inside the cave through its complex structures, and then safely return it to the surface. In addition to evaluating the geometric accuracy and robustness, we consider how informative the generated Ego-to-Exo views are compared to traditional consoles for ROV teleoperation.
@@ IV-D Subjective User Study
The user study is conducted with multiple underwater cave exploration data. A BlueROV2 recorded egocentric video feeds inside the caves at up to $100$-meter penetrations. Later on, $15$ participants evaluate the ease of operation with our developed console and compare it to traditional consoles. Their feedback is recorded using the System Usability Scale (SUS) [51], with our interface achieving an average SUS score of $77.5$. We also formulate an independent set of questions that reflect the teleoperator’s preference for the novel features of our method. The individual questions and corresponding scores are presented in Table III. Some key observations from this study are listed below.

### v3

@@ III-B Generating Exocentric Perspective
Here, $\lambda_{2}$ is the scaling factor for the ROV model. Note that our mapping and projection method is up to scale, like all monocular SLAM-based systems [47, 48]. While the scale can be resolved with additional sensor fusion, the augmented visuals of Eq. 3 are sufficient for teleoperation.
@@ IV-A Implementation Details
The framework is implemented using ROS Noetic in an Ubuntu $20.04$ environment, running on an Intel Core i9 processor with $16$ GB of RAM. A ROS node for ORB-SLAM3 is integrated as the monocular SLAM backbone. The buffer queue size $n$ is set to $100$ and the frame separation threshold is set to $0.001$ unit (up to scale). The ROV point clouds are generated by sampling 3D mesh models of BlueROV2 and TurtleBot4; $10,000$ points are sampled for each model. The scaling parameters $\lambda_{1}$ and $\lambda_{2}$ are empirically tuned once for each test sequence according to the scale of the SLAM-generated map, ensuring that the rendered ROV model appears realistic in size.
@@ IV-B Proof of Concept: 2D Indoor Navigation
Experimental setup. The proof-of-concept experiments are conducted with TurtleBot4, a 2D ground robot that can be teleoperated with egocentric views from its front-facing monocular camera. It has only two degrees of freedom (DOF) for linear and angular velocity - which simplifies the motion kinematics for tracking its instantaneous position and orientation. We teleoperate it to collect visual data with a USB camera at $640\times 480$ p resolution in office, laboratory, and hallway scenarios. The experiments are designed to validate the proposed algorithm by evaluating ground plane estimation and reprojection errors.
@@ IV-C Field Deployment: 3D Underwater Cave Exploration
Experimental setup. We extend our experiments to underwater cave exploration scenarios, where the ROV performs full 6-DOF motions. While the $roll$ motion is limited in the standard BlueROV2s, we consider all 6-DOF for teleoperation with the buoyancy change and pressure imbalance caused by water flow at the cave openings. For remote teleoperation, we consider the scenarios where human operators maneuver an underwater ROV from the surface by following the caveline and other navigation markers as a guide [49]. The mission objective is to navigate the ROV $75$-$300$ feet deep inside the cave through its complex structures, and then safely return it to the surface. The videos are recorded at $1920\times 1080$ p resolution with a GoPro11 camera mounted on BlueROV2 and then compressed to $640\times 480$ p the within the Ego-to-Exo framework. In addition to evaluating the geometric accuracy and robustness, we consider how informative the generated Ego-to-Exo views are compared to traditional consoles for ROV teleoperation.
@@ IV-D Subjective User Study
The user study is conducted with multiple underwater cave exploration data. A BlueROV2 recorded egocentric video feeds inside the caves at up to $100$-meter penetrations. Later on, $15$ human participants, between the ages of $21$-$32$ with no/little teleoperation experience, evaluate the ease of operation with our developed console and compare it to traditional consoles. Their feedback is recorded using the System Usability Scale (SUS) [51], with our interface achieving an average SUS score of $77.5$. We also formulate an independent set of questions that reflect the teleoperator’s preference for the novel features of our method. The individual questions and corresponding scores are presented in Table III. Some key observations from this study are listed below.